# Analysis of Shuffling Beyond Pure Local Differential Privacy

### v2

@@ 2.2 Local Randomizers Beyond Pure Local DP
Definition 2.6 (Generalized Gaussian local randomizer).
@@
In other words, for each input $x\in\mathcal{X}$, the output $\mathcal{R}(x)$ is obtained by adding generalized Gaussian noise (with parameters $c,\beta$) to the input $x$.
@@ 3 Analysis of Shuffling via the Central Limit Theorem (CLT)
Lemma 2.1 and Theorem 2.2 show that, for the shuffling mechanism, the DP parameter $\delta$ is bounded above and below by expressions in terms of the blanket divergence. Consequently, by analyzing the blanket divergence, we can assess the tightness of the privacy guarantees obtained via the blanket divergence.
@@ 3.1 Analysis of Blanket Divergence
Our goal here is to analyze the blanket divergence by deriving an asymptotic expansion and identifying the leading term, including the explicit constant factor. The key observation is that the blanket divergence can be expressed in terms of sums of i.i.d. random variables (even though the common distribution depends on $n$), which allows us to apply CLT-type asymptotic expansions (see Theorem A.1 and Theorem A.3 for details). To prepare for this, we first rewrite $\mathcal{D}^{\mathrm{blanket}}$ and normalize it into forms that are more amenable to the application of these theorems.
Remark 3.1 (Transformations for the blanket divergence).
The full proof is deferred to Appendix A.1. We retain both transformations because the conditions required to identify the leading term differ between the continuous (Cont) and bounded settings (Bound). In the continuous case of Assumption 2.3, we apply Theorem A.1 to the representation (1). In the bounded case (Bound), we instead apply Theorem A.3 to the representation (2). The resulting leading terms coincide, and we obtain the following result.
@@
Assume that the local randomizer $\mathcal{R}$ satisfies Assumption 2.3. Fix a pair $x_{1}\neq x_{1}^{\prime}\in\mathcal{X}$ and let $\mathcal{R}_{\mathrm{ref}}\in\bigl\{\mathcal{R}_{\mathrm{BG}}\bigr\}\,\cup\,\bigl\{\mathcal{R}_{x}:x\in\mathcal{X}\bigr\}.$ Set $\sigma^{2}:=\mathrm{Var}_{Y\sim\mathcal{R}_{\mathrm{ref}}}\bigl(l_{\varepsilon=0}(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})\bigr)$ and define $\chi:=\frac{\sqrt{\gamma}}{\sigma},$ where $\gamma\in(0,1]$ is the blanket mass of $\mathcal{R}$ when $\mathcal{R}_{\mathrm{ref}}=\mathcal{R}_{\mathrm{BG}}$ and $\gamma=1$ when $\mathcal{R}_{\mathrm{ref}}=\mathcal{R}_{x}$ for some $x$.
Assume that $\varepsilon_{n}\to 0$ as $n\to\infty$ and that $\varepsilon_{n}=\omega(n^{-1/2})$ and $\varepsilon_{n}=O\!\left(\sqrt{\frac{\log n}{n}}\right).$ Then, as $n\to\infty$,
@@
$\mathcal{D}^{\mathrm{blanket}}_{e^{\varepsilon_{n}},n,\mathcal{R}_{\mathrm{ref}},\gamma}\;=\;\varphi\bigl(\chi\bigl(\mathrm{e}^{\varepsilon_{n}}-1\bigr)\sqrt{n}\bigr)\,\left(\frac{1}{\chi^{3}}\,\frac{1}{\bigl(\mathrm{e}^{\varepsilon_{n}}-1\bigr)^{2}n^{3/2}}\right)\bigl(1+o(1)\bigr)\;=\;\varphi\bigl(\chi\varepsilon_{n}\sqrt{n}\bigr)\,\left(\frac{1}{\chi^{3}}\,\frac{1}{\varepsilon_{n}^{2}n^{3/2}}\right)\bigl(1+o(1)\bigr).$ (3)
The $o(1)$ term depends only on the constants appearing in Assumption 2.3.
See Appendix A.2 for the full proof. The leading term in (3) is obtained by replacing $F_{n-1}$ with the standard normal CDF $\Phi$ in (1) and expanding $\frac{1}{\gamma}\,\mathbb{E}\!\left[Z_{1}\,\bigl\{1-\Phi\bigl(t_{n-1}-\eta_{n-1}(Z_{1})\bigr)\bigr\}\right]$ via Mills’ ratio and a Taylor expansion. In other words, Lemma 3.2 shows that the normal approximation not only captures the correct rate, but is asymptotically exact at the level of the leading constant. The bulk of the lengthy proof is devoted to verifying that all other error contributions are asymptotically negligible compared to this main term. A crucial structural fact is that, for any local randomizer $\mathcal{R}$, one has the identity for $\mu_{\varepsilon}$ in Lemma 3.1, $\mu_{\varepsilon}=\mathbb{E}\left[Z_{1}\right]=\gamma\bigl(1-{\rm e}^{\varepsilon}\bigr),$ so that the dependence on $\varepsilon_{n}$ in the leading term of (3) takes the particularly simple form shown above.
@@
Our key conceptual finding is that the leading constant depends on the local randomizer only through the single parameter $\chi=\frac{\sqrt{\gamma}}{\sigma}.$ In particular, the relationship between the local randomizer $\mathcal{R}$ and the blanket divergence is fully encoded, at the level of the leading asymptotics, by the one-dimensional index $\chi$. Owing to its central role, we refer to $\chi$ as the shuffle index.
@@ 3.2 Main Result
Up to this point we have analyzed the blanket divergence for a fixed pair $(x_{1},x_{1}^{\prime})$ of inputs. In contrast, ($\varepsilon_{n},\delta_{n}$)-DP requires a guarantee that holds uniformly over all neighboring datasets $x_{1:n}\simeq x^{\prime}_{1:n}$ (i.e., $\delta_{n}=\sup_{x_{1:n}\simeq x^{\prime}_{1:n}}\mathcal{D}_{e^{\varepsilon_{n}}}(\mathcal{M}(x_{1:n})\|\mathcal{M}(x^{\prime}_{1:n}))$). Our next goal is therefore to extend the fixed-pair asymptotics of Lemma 3.2 to an asymptotically worst-case guarantee over all neighboring datasets.
Recall that Lemma 3.2 identifies, for each fixed pair $(x_{1},x_{1}^{\prime})$ and each choice of reference distribution $\mathcal{R}_{\mathrm{ref}}$, a shuffle index $\chi(x_{1},x_{1}^{\prime};\mathcal{R}_{\mathrm{ref}})$ such that the corresponding blanket divergence has the asymptotic form (3). In particular, the dependence on the local randomizer enters the leading term only through the single scalar $\chi$. Consequently, the problem of upper bounding $\delta_{n}$ reduces to understanding the worst value of $\chi$ over all pairs $(x_{1},x_{1}^{\prime})$ under $\mathcal{R}_{\mathrm{ref}}=\mathcal{R}_{\mathrm{BG}}$. Establishing the tightest lower bound on $\delta_{n}$ corresponds to finding the pair $(x_{1},x_{1}^{\prime})$ and the reference distribution $\mathcal{R}_{\mathrm{ref}}\in\{\mathcal{R}_{x}:x\in\mathcal{X}\}$ that minimize the shuffle index. We now state the result formally.
@@
See Appendix A.3 for the full proof. Informally, Theorem 3.3 shows that for any fixed target constant $\alpha>0$ we construct a sequence $(\varepsilon_{n}^{\ast})_{n\geq 2}$ whose values lie between two explicit curves $\varepsilon_{n}(\alpha,\underline{\chi}_{\mathrm{up}})$ and $\varepsilon_{n}(\alpha,\underline{\chi}_{\mathrm{lo}})$, and for which $\delta_{n}=\frac{\alpha}{n}(1+o(1))$. [^4] More generally, one can obtain any prescribed polynomial rate $\delta_{n}=\Theta(n^{-\beta})$ with $\beta>1$ in the leading term. In this work we focus on the regime $\delta_{n}=\alpha/n(1+o(1))$ because in the DP literature, setting $\delta_{n}$ to be $\alpha/n$ is a common practice. In other words, the band between $\varepsilon_{n}(\alpha,\underline{\chi}_{\mathrm{up}})$ and $\varepsilon_{n}(\alpha,\underline{\chi}_{\mathrm{lo}})$ precisely captures $\varepsilon^{\ast}_{n}$ that realizes $\delta_{n}\approx\alpha/n$.

### v3

@@ 2.2 Local Randomizers Beyond Pure Local DP
We denote by $\mathfrak{R}$ the class of local randomizers $\mathcal{R}$ that satisfy Assumption 2.3.
@@
Definition 2.6 (Generalized Gaussian local randomizer ($\beta$-Gaussian mechanism)).
@@
In other words, for each input $x\in\mathcal{X}$, the output $\mathcal{R}(x)$ is obtained by adding generalized Gaussian noise (with parameters $c,\beta$) to the input $x$. The variance of this distribution is given by $\mathrm{Var}(Y)=c^{2}\,\frac{\Gamma(3/\beta)}{\Gamma(1/\beta)}$; by choosing $c$ appropriately, we can set $\mathrm{Var}(Y)=\sigma_{0}^{2}.$ We call $\mathcal{R}$ the $\beta$-Gaussian mechanism.
@@ 3 Analysis of Shuffling via the Central Limit Theorem (CLT)
Here, we first review the blanket divergence and its interpretation. Then, we analyze the blanket divergence to obtain an asymptotic expansion. Next, we analyze the privacy profile of the shuffled mechanism based on this expansion. Then, we interpret and explore the shuffle indices of various local randomizers. Finally, we derive a necessary and sufficient condition under which the blanket divergence yields an asymptotically optimal characterization.
3.1 Review of Blanket Divergence
Given a shuffled mechanism $\mathcal{M}=\mathcal{S}\circ\mathcal{R}^{n}$, the direct computation of the privacy profile requires the two output distributions of $\mathcal{M}(x_{1:n})$ and $\mathcal{M}(x^{\prime}_{1:n})$ where $x_{1:n}$ and $x^{\prime}_{1:n}$ are the worst-case neighboring datasets. Given a dataset $x_{1:n}$, the output distribution of the shuffled mechanism $\mathcal{M}=\mathcal{S}\circ\mathcal{R}^{n}$ can be written as the permutation-mixture
$\Pr[\mathcal{M}(x_{1:n})=y_{1:n}]=\frac{1}{n!}\sum_{\pi\in S_{n}}\;\prod_{i=1}^{n}\mathcal{R}_{x_{i}}\bigl(y_{\pi(i)}\bigr),\qquad y_{1:n}\in\mathcal{Y}^{n}.$
That is, evaluating the shuffled density naively involves a mixture with $n!$ components. Moreover, we need to apply a nonlinear transformation to this and integrate the result, together with a maximization over all neighboring datasets to identify the worst case. Altogether, this direct computation quickly becomes intractable. We therefore turn to the privacy blanket approach proposed by Balle et al. (2019).
Here, we briefly recall it. The local randomizer admits the mixture decomposition $\mathcal{R}_{x}=\gamma\,\mathcal{R}_{\mathrm{BG}}+(1-\gamma)\,\mathcal{R}_{x}^{\mathrm{LO}},$ where $\mathcal{R}_{x}^{\mathrm{LO}}$ is the corresponding leftover distribution. Based on this decomposition, we consider the extended mechanism $\widetilde{\mathcal{M}}$ depicted in Fig. 1. We can interpret a blanket divergence as an upper bound on the hockey-stick divergence of $\widetilde{\mathcal{M}}$.
In $\widetilde{\mathcal{M}}$, each user ($\mathrm{id}>1$) draws a message from $\mathcal{R}_{\mathrm{BG}}$ with probability $\gamma$, and from $\mathcal{R}_{x_{\mathrm{id}}}^{\mathrm{LO}}$ with probability $1-\gamma$. Messages sampled from the leftover part are revealed together with their user indices, whereas messages sampled from the blanket part are released only after being shuffled. User $1$ outputs $\mathcal{R}_{x_{1}}$ (resp. $\mathcal{R}_{x_{1}^{\prime}}$) and $Y_{1}$ is always placed in the shuffled blanket multiset. The hockey-stick divergence of $\widetilde{\mathcal{M}}$ is upper-bounded by applying its joint convexity to the mixture over the number of blanket messages, which yields the blanket divergence (see Section B for details). Importantly, the shuffled mechanism $\mathcal{M}=\mathcal{S}\circ\mathcal{R}^{n}$ can be obtained by shuffling all outputs of $\widetilde{\mathcal{M}}$. Therefore, by the data processing inequality, the hockey-stick divergence of $\mathcal{M}$ is also upper-bounded by the blanket divergence.
Next, in Fig. 1, we consider a fixed $x_{2}\in\mathcal{X}$ and restrict attention to the neighboring-dataset instance $(x_{1},x_{2},\dots,x_{2})\ \simeq\ (x_{1}^{\prime},x_{2},\dots,x_{2}).$ In this setting, if we take $\gamma=1$ and choose the reference distribution as $\mathcal{R}_{\mathrm{BG}}=\mathcal{R}_{x_{2}}$, then the resulting construction of $\widetilde{\mathcal{M}}$ coincides in distribution with the shuffled mechanism output for these two datasets (i.e., $\mathcal{M}((x_{1},x_{2},\dots,x_{2}))$ and $\mathcal{M}((x_{1}^{\prime},x_{2},\dots,x_{2}))$). The inequalities introduced earlier, namely the convexity step in $\gamma$ and the data processing inequality, do not appear here, so that the blanket divergence matches the hockey-stick divergence of $\widetilde{\mathcal{M}}$, and thus that of the shuffled mechanism $\mathcal{M}$. Since this neighboring instance is feasible in the supremum that defines the privacy profile, it follows that this blanket divergence provides a lower bound.
Remark 3.1 (Relationship to the clone paradigm).
The blanket divergence viewpoint with $\gamma=1$ is closely related to the clone paradigm Feldman et al. (2022, 2023). The clone paradigm reduces shuffle privacy analysis to an auxiliary mechanism; shuffling of $n$ categorical distributions $(\mathcal{R}^{\mathrm{Cat}}_{b},\mathcal{R}^{\mathrm{Cat}}_{\mathrm{ref}},\dots,\mathcal{R}^{\mathrm{Cat}}_{\mathrm{ref}})$, where $b\in\{0,1\}$ changes the target user. In this sense, this reduced mechanism can be interpreted as a particular $\gamma=1$ instance of $\widetilde{\mathcal{M}}$, with an appropriate choice of $\mathcal{R}^{\mathrm{Cat}}_{b}$ and $\mathcal{R}^{\mathrm{Cat}}_{\mathrm{ref}}$. Note that Su et al. (2025) showed that the blanket-based reduction can be at least as tight as the clone reductions. Motivated by this and by the mechanism-aware nature of the blanket divergence, we focus primarily on the blanket-divergence upper bound of Balle et al. (2019) in what follows.
Figure 1: $\widetilde{\mathcal{M}}$, which yields the blanket divergence.
Figure 2: The intuition behind the CLT of the blanket divergence with $m=100$.
Figure 3: The intuition of $\sigma$ with the Gaussian mechanism where $\sigma_{0}=2$. $x$ axis represents the output of the local randomizer and is common for both images.
@@ 3.2 Analysis of Blanket Divergence
As shown in Lemma 2.1, the blanket divergence no longer involves an explicit permutation. This is because, conditioned on the number of blanket messages $M=m$ in $\widetilde{\mathcal{M}}$, the m1m-1 blanket outputs are i.i.d. draws from the blanket distribution $\mathcal{R}_{\mathrm{BG}}$; moreover, thanks to the symmetry attained by shuffling, the analysis of the hockey-stick divergence reduces to that of a sum of $m$ i.i.d. random variables. Exploiting this tractability via the CLT, we can derive the following asymptotic expansion.
@@
Given $\mathcal{R}\in\mathfrak{R}$, fix a pair $x_{1}\neq x_{1}^{\prime}\in\mathcal{X}$ and let $\mathcal{R}_{\mathrm{ref}}\in\bigl\{\mathcal{R}_{\mathrm{BG}}\bigr\}\,\cup\,\bigl\{\mathcal{R}_{x}:x\in\mathcal{X}\bigr\}.$ For $\gamma\in(0,1]$, define $\sigma^{2}:=\mathrm{Var}_{Y\sim\mathcal{R}_{\mathrm{ref}}}\bigl(l_{\varepsilon=0}(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})\bigr)$ and $\chi:=\sqrt{\gamma}/{\sigma}.$ Assume that as $n\to\infty$, $\varepsilon_{n}=\omega(\sqrt{1/n})$ and $\varepsilon_{n}=O\!\left(\sqrt{\log n/n}\right).$ Then, as $n\to\infty$,
@@
$\mathcal{D}^{\mathrm{blanket}}_{e^{\varepsilon_{n}},n,\mathcal{R}_{\mathrm{ref}},\gamma}(\mathcal{R}_{x_{1}}\|\mathcal{R}_{x_{1}^{\prime}})$ $\;=\;\varphi\bigl(\chi\varepsilon_{n}\sqrt{n}\bigr)\,\left(\frac{1}{\chi^{3}}\,\frac{1}{\varepsilon_{n}^{2}n^{3/2}}\right)\bigl(1+o(1)\bigr).$ (1)
Here, we present an informal derivation based on the CLT to build intuition of how the result arises. The rigorous proof with the asymptotic expansion of CLT Slastnikov (1979); Angst and Poly (2017); Petrov (1975) is in Appendix A.1.
Recall that the blanket divergence can be written as
$\mathcal{D}^{\mathrm{blanket}}=\frac{1}{n\gamma}\;\mathbb{E}_{M}\!\left[\mathbb{E}\!\left[\Bigl(\sum_{i=1}^{M}l_{\varepsilon_{n}}(Y_{i})\Bigr)_{+}\,\Big|\,M\right]\right],\qquad M\sim\mathrm{Bin}(n,\gamma),\;\;Y_{i}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{R}_{\mathrm{ref}}.$
Let us condition on $M=m$ and define $S_{m}:=\sum_{i=1}^{m}l_{\varepsilon_{n}}(Y_{i}).$ The inner expectation becomes $\mathbb{E}[(S_{m})_{+}]$, whose behavior is shown in Fig. 2. Since $S_{m}$ is a sum of i.i.d. random variables, the CLT suggests
$S_{m}\approx N_{m}:=\mathcal{N}(\mu_{m},s_{m}^{2}),\qquad\mu_{m}:=m\,\mathbb{E}[l_{\varepsilon_{n}}(Y)]=m(1-e^{\varepsilon_{n}}),\quad s_{m}^{2}:=m\,\mathrm{Var}(l_{\varepsilon_{n}}(Y)),$
for large $m$. For a Gaussian $N_{m}\sim\mathcal{N}(\mu_{m},s^{2}_{m})$ one has the identity $\mathbb{E}[(N_{m})_{+}]=s_{m}\,\varphi\!\left(\frac{\mu_{m}}{s_{m}}\right)+\mu_{m}\,\Phi\!\left(\frac{\mu_{m}}{s_{m}}\right).$ Using Mills’ ratio (in the regime $|\mu_{m}/s_{m}|\to\infty$), we obtain
$\mathbb{E}[(N_{m})_{+}]\;\approx\;\frac{s_{m}^{3}}{\mu_{m}^{2}}\,\varphi\!\left(\frac{\mu_{m}}{s_{m}}\right)\;=\;\frac{\mathrm{Var}(l_{\varepsilon_{n}}(Y))^{3/2}}{(e^{\varepsilon_{n}}-1)^{2}}\,\frac{1}{\sqrt{m}}\,\varphi\!\left(\frac{(e^{\varepsilon_{n}}-1)\sqrt{m}}{\sqrt{\mathrm{Var}(l_{\varepsilon_{n}}(Y))}}\right),$
$M\sim\mathrm{Bin}(n,\gamma)$ concentrates around its mean $m=n\gamma$, so we plug in $m=n\gamma$ in the leading term. Moreover, for small $\varepsilon_{n}$, $\mathrm{Var}(l_{\varepsilon_{n}}(Y))\approx\mathrm{Var}(l_{0}(Y))=:\sigma^{2}$. With $\chi:=\sqrt{\gamma}/\sigma$, we obtain
$\mathbb{E}_{M}[(N_{M})_{+}]\approx\frac{\sigma^{3}}{(e^{\varepsilon_{n}}-1)^{2}}\;\frac{1}{\sqrt{n\gamma}}\;\varphi\!\bigl(\chi(e^{\varepsilon_{n}}-1)\sqrt{n}\bigr).$
Finally, recalling the prefactor $1/(n\gamma)$, and using $\sigma^{3}/\gamma^{3/2}=1/\chi^{3}$, we arrive at
$\mathcal{D}^{\mathrm{blanket}}\approx\frac{1}{n\gamma}\,\mathbb{E}_{M}[(N_{M})_{+}]\approx\varphi\!\bigl(\chi(e^{\varepsilon_{n}}-1)\sqrt{n}\bigr)\,\left(\frac{1}{\chi^{3}}\,\frac{1}{(e^{\varepsilon_{n}}-1)^{2}\,n^{3/2}}\right),$
which matches the leading term in (1) using $\mathrm{e}^{\varepsilon_{n}}\approx 1+\varepsilon_{n}$. The rigorous proof in Appendix A.1 shows that this normal approximation is indeed valid under Assumption 2.3 (i.e., $\mathcal{R}\in\mathfrak{R}$) in the regime $\varepsilon_{n}=\omega(\sqrt{1/n})$ and $\varepsilon_{n}=O(\sqrt{\log n/n})$.
@@
Our key conceptual finding from Lemma 3.2 is that the leading term depends on the local randomizer only through the single parameter $\chi=\sqrt{\gamma}/{\sigma}.$ That is, at the level of the leading asymptotics, the dependence of the blanket divergence on the local randomizer $\mathcal{R}$ is fully encoded by the one-dimensional index $\chi$. Furthermore, the leading term is monotone in $\chi$: a larger $\chi$ yields a smaller leading term in the blanket divergence, and hence corresponds to stronger privacy amplification by shuffling. This motivates interpreting $\chi$ as a measure of the shuffle efficiency of the local randomizer. Owing to its central role, we refer to $\chi$ as the shuffle index.
@@ 3.3 Main Result
Up to this point, we have analyzed the blanket divergence for a fixed pair $(x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})$. In contrast, ($\varepsilon_{n},\delta_{n}$)-DP requires a guarantee that holds uniformly over all neighboring datasets $x_{1:n}\simeq x^{\prime}_{1:n}$ (i.e., upper bounding the privacy profile $\delta_{n}\geq\delta_{\mathcal{M}}(\varepsilon_{n})$). This section aims to relate the result of the blanket divergence to the privacy profile of the shuffled mechanism.
Recall from Section 3.1 that the privacy profile can be sandwiched between two blanket divergences. By combining this property with the monotonicity of the leading term with respect to the shuffle index $\chi$, evaluating the privacy profile reduces to identifying the worst-case value of $\chi$ over all pairs $(x_{1},x_{1}^{\prime})$. To derive the upper bound, under the reference distribution $\mathcal{R}_{\mathrm{ref}}=\mathcal{R}_{\mathrm{BG}}$, we must choose the pair $(x_{1},x_{1}^{\prime})$ that minimizes the shuffle index. Conversely, to establish the lower bound, under $\mathcal{R}_{\mathrm{ref}}=\mathcal{R}_{x}$ for some $x\in\mathcal{X}$ (with $\gamma=1$), we are free to select any pair $(x_{1},x_{1}^{\prime})$; naturally, to maximize this lower bound, we should again choose the pair that minimizes the shuffle index. Formally, we use the following specific shuffle indices for the upper and lower bounds, respectively.
Definition 3.1.
Let $\gamma\in(0,1]$ be the blanket mass of $\mathcal{R}\in\mathfrak{R}$. The lower shuffle index is defined by $\chi_{\mathrm{lo}}(\mathcal{R}):=\sqrt{\gamma}/\sup_{x_{1}\simeq x_{1}^{\prime}\in\mathcal{X}}\sqrt{\mathrm{Var}_{Y\sim\mathcal{R}_{\mathrm{BG}}}\!\left[l_{0}\!\left(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{BG}}\right)\right]}.$ The upper shuffle index is defined by $\chi_{\mathrm{up}}(\mathcal{R}):=1/\sup_{x_{1}\simeq x_{1}^{\prime}\in\mathcal{X}}\ \sup_{x\in\mathcal{X}}\sqrt{\mathrm{Var}_{Y\sim\mathcal{R}_{x}}\!\left[l_{0}\!\left(Y;x_{1},x_{1}^{\prime},\mathcal{R}_{x}\right)\right]}.$
With these specific shuffle indices in hand, we can formally establish the upper and lower bounds for the privacy profile of the shuffled mechanism, as stated in the following corollary.
Remark 3.3 (Privacy profile bounds).
Let $\mathcal{R}\in\mathfrak{R}$. Assume that as $n\to\infty$, $\varepsilon_{n}=\omega(\sqrt{1/n})$ and $\varepsilon_{n}=O\!\left(\sqrt{\log n/n}\right).$ Then, there exist sequences $e_{n}^{\mathrm{up}},e_{n}^{\mathrm{lo}}\to 0$ as $n\to\infty$ such that
$f_{n,\varepsilon_{n}}(\chi_{\mathrm{up}}(\mathcal{R}))\,\bigl(1+e_{n}^{\mathrm{up}}\bigr)\leq\delta_{\mathcal{S}\circ\mathcal{R}^{n}}(\varepsilon_{n})\leq f_{n,\varepsilon_{n}}(\chi_{\mathrm{lo}}(\mathcal{R}))\,\bigl(1+e_{n}^{\mathrm{lo}}\bigr),$
for all sufficiently large $n$ where $f_{n,\varepsilon_{n}}(\chi):=\varphi\bigl(\chi\,\varepsilon_{n}\sqrt{n}\bigr)\,\left(\frac{1}{\chi^{3}}\,\frac{1}{\varepsilon_{n}^{2}n^{3/2}}\right).$
This corollary is a direct consequence of applying the asymptotic expansion established in Lemma 3.2 to the lower and upper bounds (Lemma 2.1 and Theorem 2.2).
In DP, it is common to set $\delta_{n}$ to be $\alpha/n$ for some fixed constant $\alpha>0$. Our next result characterizes the corresponding $\varepsilon_{n}$.
@@
$\varepsilon_{n}(\alpha,\chi):=\log\left(1+\sqrt{\frac{2}{\chi^{2}n}\,W\!\left(\frac{\sqrt{n}}{2\alpha\chi\sqrt{2\pi}}\right)}\right)$ (2)
Then, there exists a sequence $(\varepsilon^{\ast}_{n}\in[\varepsilon_{n}\bigl(\alpha,\chi_{\mathrm{up}}(\mathcal{R})\bigr),\varepsilon_{n}\bigl(\alpha,\chi_{\mathrm{lo}}(\mathcal{R})\bigr)])_{n\geq 1}$ such that $\delta_{\mathcal{S}\circ\mathcal{R}^{n}}(\varepsilon^{\ast}_{n})=\frac{\alpha}{n}(1+o(1))$ for all sufficiently large $n$.
See Appendix A.2 for the full proof. This is obtained by inverting the relationship between $\varepsilon_{n}$ and $\delta_{n}$ from the bounds established in Corollary 3.3. Specifically, by equating the leading term of the privacy profile $f_{n,\varepsilon_{n}}(\chi)$ to the target $\delta_{n}=\alpha/n$, we can solve for $\varepsilon_{n}$ to derive the explicit formula $\varepsilon_{n}(\alpha,\chi)$ for both the upper and lower bounds. Informally, Theorem 3.4 shows that for any fixed target constant $\alpha>0$ we construct a sequence $(\varepsilon_{n}^{\ast})_{n\geq 2}$ whose values lie between two explicit curves $\varepsilon_{n}(\alpha,\chi_{\mathrm{up}}(\mathcal{R}))$ and $\varepsilon_{n}(\alpha,\chi_{\mathrm{lo}}(\mathcal{R}))$, and for which $\delta_{n}=\frac{\alpha}{n}(1+o(1))$. [^4] More generally, one can obtain any prescribed polynomial rate $\delta_{n}=\Theta(n^{-\beta})$ with $\beta>1$ in the leading term. In this work we focus on the regime $\delta_{n}=\alpha/n(1+o(1))$ because in the DP literature, setting $\delta_{n}$ to be $\alpha/n$ is a common practice. In other words, the band precisely captures $\varepsilon^{\ast}_{n}$ that realizes $\delta_{n}\approx\alpha/n$.
[…]
$\frac{\varepsilon_{n}(\alpha,\chi_{\mathrm{lo}}(\mathcal{R}))}{\varepsilon_{n}(\alpha,\chi_{\mathrm{up}}(\mathcal{R}))}\;=\;\frac{\chi_{\mathrm{lo}}(\mathcal{R})}{\chi_{\mathrm{up}}(\mathcal{R})}\,\left(1+o(1)\right).$
Hence, if $\chi_{\mathrm{lo}}(\mathcal{R})/\chi_{\mathrm{up}}(\mathcal{R})$ is close to 1, the band is narrow, indicating that the shuffle index provides a tight characterization of the amplified privacy parameter.